Find the distance between the two points (-4, -5) and (1, -2).
Answer by richwmiller(9135)

c^2=x^2+y^2
y=(-5)-(-2)=-5+2=-3
x=-4-(+1)=-4-1=-5
c^2=(-4)^2+(-5)^2
c^2=16+25=41
c=Sqrt(41)=abt 6.4